if all the students with an iq of 87 and lower will be required to attend the remedial program, how many are they?



1
Expert's answer
2021-04-29T17:08:37-0400

The office for admission of XYZ University found out that the IQ of incoming 1200 freshmen normally

distributed with a mean of 108 and a standard deviation of 10.

So, according to question:

Mean "\mu=108"

Standard Deviation "\sigma=10"

n= 1200

"P(X<87)=\dfrac{x-\mu}{\sigma}=\dfrac{87-108}{10}=-2.1"





P(z<-2.1)=0.0179

Hence, No. of students  required to attend the remedial program are = 0.0179 "\times" 1200=21.48

So, approx. 22 students are of IQ of 87 and lower than are required to attend the remedial program.